Fermi National Laboratory


Accelerator Update for December 17, 2004 - December 20, 2004

Friday December 17
The day shift began with Operations monitoring store 3853, with the Antiproton Source (Pbar) stacking, with the Recycler conducting studies, and with beam to MiniBooNE and the Meson experiment MTest.

At 3:11 PM, a MiniBooNE power supply (HV860) suffered a LCW leak that held off beam for an hour.

At 10:39 PM, Operations reported that Booster tripped off and vacuum valves closed.

At 10:45 PM, Operations reported that the Tevatron (TeV) sector F3 area suffered an air pressure drop for vacuum valves. The air compressor came on and the system recovered.

At 10:47 PM, Operations reported an emittance blowup in Pbar, which was causing antiprotons to unstack. Operators called an expert for help.

Operators had the Booster trip reset and had all valves open by 11 PM.

Operations reported by 11:45 PM that Pbar had lost about 4mA. Operators tried several fixes with little or no help. They played with the horizontal dampers and that helped a little bit. (The blowup occurred just before the Booster trip.)

At 11:47 PM, a Booster sump went into alarm and would not clear.

Saturday December 18
The midnight shift began with Operations monitoring store 3853, with Pbar stacking, and with MiniBooNE and MTest taking beam.

Operations began preparing Booster for access at 12:10 AM to investigate the sump problem. Operators found no problem with the sump.

Operations had re-established beam to MI and to MiniBooNE by 1:18 AM. They held off the Pbar beam until an expert arrived and helped with the emittance problem. Stacking resumed around 1:43 AM.

Operations terminated store 3853 at 3:03 AM and put the TeV into shot setup at 3:14 AM.

Operations established store 3854 at 5:50 AM with an initial luminosity 61.06E30.

At 6:05 AM, the TeV suffered a sector E1 quench. Operators investigated and found that a sector B1 separator spark had caused the beam to blowup and hit the D4 collimator, which caused the quench.

Operations put the TeV separators into conditioning mode at 6:41 AM, per TeV expert suggestion.

Operators began preparing the D-Zero collision hall for an access at 7:04 AM and prepared the Collider Detector at Fermilab (CDF) for an access at 9:32 AM.

A TeV expert began an octupole study at 9:50 AM.

CDF came out of access at 10:33 AM.

At 1:18 PM, the 400 MeV area suffered the first of four trips over a five-hour period.

At 4:08 PM, the TeV suffered a sector F1 quench. The cause was due to injecting beam with two separators on.

A TeV expert began a study period at 6:02 PM.

At 9:12 PM, a MiniBooNE power supply was reading hot. Operators investigated but didn't see a problem. They called in an expert. This held MiniBooNE off.

Operations put the TeV into shot setup at 9:55 PM.

An expert brought the MiniBooNE power supply back on line at 10:33 PM.

Sunday December 19
The midnight shift began with the TeV in shot setup, and with MiniBooNE and MTest taking beam.

Everything had been loaded for the next store when the TeV suffered a quench during a low beta squeeze at 12:50 AM. Operators and a TeV expert began investigated and initially thought a quadrupole (DQ7) had pulled the abort, but experts later decided this wasn't the cause of the quench. Investigations continued. Pbar resumed stacking.

Operations established beam to MTest at 1:51 AM.

At 1:59 AM, a Booster RF station (BRF2) tripped off. Operators investigated and tried several resets, but nothing worked. Operators bypassed the station.

At 2:47 AM, the H- Source's output dropped to a very low output. A Linac expert came in to investigate at 4:11 AM.

The Linac expert switched over to the I- Source at 5:39 AM.

A TeV expert began studies at 7:37 AM.

At 9:14 AM, a TeV sector cryo system began oscillating. Cryo technicians from CHL investigated and found a bad control card for the cold compressor. The techs had the system back to normal about two hours later.

At 9:47 AM, a Linac RF station (LRF1) tripped off and wouldn't reset. Operators called in experts.

Linac experts began repairs on LRF1 at 1:27 PM. Among other problems, the techs found a high voltage cable that had burned open.

high voltage cable that had burned open

LRF1 repairs finished at 4:24 PM. Pbar resumed stacking at 4:45 PM.

At 5:15 PM, Operations reported the 400 MeV area had tripped off for the first of four times in seven minutes.

At 6:35 PM, the Pbar lithium lens tripped off. It reset okay.

Operations put the TeV into shot setup at 6:42 PM.

At 7:09 PM, Operations reported an A49 separator spark.

The Pbar beamline tune up for shot setup had a problem at 9:03 PM. Operators contacted Pbar experts.

At 9:48 PM, Operations reported that they had seen four sparks registered for the A49 separator, but they'd seen no affect on the beam.

Operations established store 3862 at 11:06 PM with an initial luminosity of 40.33E30.

At 11:14 PM, Operations reported a TeV A17 separator spark.

At 11:25 PM, the TeV store aborted due to a problem with a quadrupole (B0Q3).

Operations re-established beam to MTest at 11:38 PM.

Pbar resumed stacking at 11:43 PM.

Monday December 20
The midnight shift began with Operations investigating the TeV abort, with Pbar stacking, and with MiniBooNE, MCenter, and MTest taking beam.

Operations put the TeV into conditioning mode at 12:10 AM.

A TeV expert arrived at 1:21 AM to help with the TeV investigations.

At 3:16 AM, Operations reported that the Pbar lithium lens tripped off. Operators reset the lens and had Pbar stacking by 3:41 AM.

At 4:10, the lithium lens tripped off again. An Pbar expert investigated with the operators and they had Pbar stacking by 4:52 AM.

A Recycler expert began at study at 5 AM.

TeV experts began a study period at 5:02 AM.

Operations took the TeV separators out of conditioning mode at 5:47 AM.

At 6:39 AM, the lithium lens tripped off. It tripped again at 7:10 AM and then again at 7:21. Operators called in an expert.

Operations put Pbar back into stacking mode.

Plans for Monday
The plan for today is to work on getting the TeV into a stable run mode by conducting several short low beta runs and then conducting a quadrupole study.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/20/2004   email Fermilab